OpenStudy (anonymous):

Please help! Really hard! ): Given F(x)=4x^3 and G(x)=2x what is F(2) divided by G(2)

OpenStudy (anonymous):

A. 2 B. 4 C. 8 D. 16

OpenStudy (mandre):

\[\frac{ F(x) }{ G(x) }=\frac{ 4x ^{3} }{ 2x }\] Plug in 2 where x is and simplify.

OpenStudy (anonymous):

oh okay when i simplified i got 4 is this correct? @Mandre

OpenStudy (mandre):

Nope. What is 2^3?

OpenStudy (anonymous):

8 (:

OpenStudy (mandre):

So? Can you see where you went wrong?

OpenStudy (anonymous):

yes i do! thanks alot! would you be able to help me with this question? Given F(x)=5x^6 and G(x)=15x^-2 what is F(x) divided by G(x)

OpenStudy (mandre):

Do you have a value for x?

OpenStudy (anonymous):

no not in this one

OpenStudy (mandre):

Did you post that as a separate question?

OpenStudy (mandre):

I thought I saw it there.

OpenStudy (anonymous):

yeah i think but know one helped me

OpenStudy (mandre):

Your equation is Given F(x)=5x^6 and G(x)=15x^-2 what is F(x) divided by G(x) \[\frac{ 5x ^{6} }{ 15x ^{-2} }\] Simplifies to \[\frac{ 5x ^{6} }{ 15/x ^{2} }\] That simplifies to \[5x ^{6}(\frac{ x ^{2} }{ 15 })\]

OpenStudy (mandre):

You can take it from there?

OpenStudy (anonymous):

No i have no idea what the answer is );

OpenStudy (mandre):

Just multiply the 5x^6 into the fraction and simplify.

OpenStudy (mandre):

\[\frac{ 5x ^{6}.x ^{2} }{ 15 }\] Surely you can simplify that?

OpenStudy (anonymous):

would the answer be 3x^8?

OpenStudy (mandre):

You have the 3 in the wrong place. Look at my last equation.

OpenStudy (anonymous):

im so confused is it x^8/3? @Mandre

OpenStudy (mandre):

There you go :)

OpenStudy (anonymous):

oh haha thank you!

OpenStudy (mandre):

Pleasure.